OpenStudy (anonymous):

Solve for x: 1/2 + 4/2x=x+4/10

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

1/2-2/5=x-2/x x/10=x^2-2 10x^2-x-20=0 then simplify get x value.

OpenStudy (goten77):

^lol?

OpenStudy (anonymous):

lol simplify only